OpenStudy (anonymous):

How are the different types of linear equations similar and how are they different?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

I'd suggest you begin by finding these three types of equations of straight lines and typing them out here. That'll give you something concrete to compare and contrast. Hint: One of the types of equations of straight lines is the slope-intercept form. What are the names of the other types?

OpenStudy (anonymous):

The slope-intercept form: y = m x + b. The point-slope form: y − y1 = m (x − x1). The general form: a x + b y + c = 0.